The agency felt pretty good with your findings in Q2 but did have qualms about using some letters and numbers that can be easily confused. Following the same rules as in Q2, but adding a rule that the capital letters I and L and number 1 can't be used, find how many possible outcomes exist.

Acceptable license plate: 892-KKT

Unacceptable license plate: 123-IJL

Group of answer choices

5,355,504

11,390,625

6,120,576

6,967,296

Q2 : The agency is also considering having a license plate with six slots where the first three slots have to be a number 0-9 (with no duplicate numbers allowed), and the last three slots must be an upper-case letter (with duplicate letters allowed). An example of a valid license plate under this rule is 423-PBN.
Answer #1

Out of the 6 places, first three cannot have 1. Hence the first three digits can be filled only by 9 numbers in 9C3*3! ways

The second 3 places can be filled only by 24 letters but can be repeated. Hence, number of ways=24*24*24=24^3 Ways

Hence total number of ways=3!*9C_3*24^3 =6967296Ways